OpenStudy (anonymous):

derivative of root 3 will be 0, right?

OpenStudy (anonymous):

ys

OpenStudy (anonymous):

then solve this: \[(x ^{3}\div \sqrt{3}) - \sqrt{3}x ^{3}\]

OpenStudy (amistre64):

root 3 is a constant; so yes ;)

OpenStudy (anonymous):

then differentiate the above equation for me, please? because that's what I am doing and the answer doesn't seem to match. will be apply the quotient or product rule? or just differentiate it directly?

OpenStudy (bahrom7893):

x^3/sqrt(3) - sqrt(3) * x^3, the derivative is: 3x^2/sqrt(3) - 3sqrt(3)*x^2

OpenStudy (amistre64):

\[Dx[\frac{x^3}{\sqrt{3}}-\sqrt{3}x^3]=Dx[\frac{x^3}{\sqrt{3}}]-Dx[\sqrt{3}x^3]\]

OpenStudy (amistre64):

\[\frac{3x^2}{\sqrt{3}}-3\sqrt{3}x^2\] is what I get from it

OpenStudy (amistre64):

its just the power rule applied after pulling out the constants; then multiply the constants back in

OpenStudy (anonymous):

what about the first term? shouldn't we not apply the quotient rule?

OpenStudy (amistre64):

\(\frac{1}{\sqrt{3}}x^3\) doesnt require the quotient rule any more than any other constant; since the variable is not a part of the bottom

OpenStudy (amistre64):

the end result, if I see it right is: -2 sqrt(3) x^2

OpenStudy (amistre64):

if you wanna DO the quotient rule; it works out the same: \[\frac{3\sqrt{3}\ x^2-0(x^3)}{3}\]

OpenStudy (anonymous):

oh okay, thanks a lot :)

OpenStudy (amistre64):

youre welcome :)